OpenStudy (anonymous):

Differentiate y=e^square root (e^2x) sorry bout the "square root" I'm using iPad please show steps for y prime

OpenStudy (jamesj):

Do you mean this: \[ y = \exp(\sqrt{e^{2x}}) \]

OpenStudy (anonymous):

I mean y=e^everything inside your brackets

OpenStudy (jamesj):

Right. So use the chain rule. For example let u(x) = e^2x, then your function is \[ y = e^{\sqrt{u(x)}} \]

OpenStudy (jamesj):

In other words, \[ \frac{dy}{dx} = \frac{dy}{du} \frac{du}{dx} \] Now calculate these two terms on the right.

OpenStudy (jamesj):

following?

OpenStudy (anonymous):

Yeah thanks sorry trying to figure out if I need to use chain rule twice ?

OpenStudy (jamesj):

Yes, you do.

OpenStudy (anonymous):

Ahhh yeah I get ya :-) thanks heaps for that james

OpenStudy (ganpat):

it might be, y = \[y = e ^\sqrt([2x]) * -2\sqrt{x}\]

OpenStudy (ganpat):

* = multiplication..

OpenStudy (jamesj):

\[ \frac{dy}{du} = \frac{1}{2\sqrt{u}} e^{\sqrt{u}} \]

OpenStudy (ganpat):

where come division comes from ?

OpenStudy (jamesj):

because the derivative of \( \sqrt{u} \) is \( 1/(2\sqrt{u}) \).

OpenStudy (ganpat):

ahhh. k k, i got confuse.. u seems right..